OpenStudy (anonymous):

simply the following limit algebraically limit as x-> 0 (x^2)/(1-cos^2(x))

OpenStudy (blockcolder):

\[1-\cos^2x=\sin^2x\] So your limit is reduced to: \[\lim_{x \to 0} \frac{x^2}{\sin^2x}=\lim_{x \to 0} \left ( \frac{x}{\sin x}\right )^2=\left ( \lim_{x \to 0} \frac{x}{\sin x}\right )^2=1\] because the squaring function is continuous.